# Accurately simulating core-collapse self-interacting dark matter halos

### v2

@@ Abstract
Context. The properties of satellite halos provide a promising probe for dark matter (DM) physics. Observations motivate current efforts to explain surprisingly compact DM halos. If DM is not collisionless but has strong self-interactions, halos can undergo gravothermal collapse, leading to higher densities in the central region of the halo. However, it is challenging to model this collapse phase from first principles.
Aims. To improve on this, we seek to better understand numerical challenges and convergence properties of self-interacting dark matter (SIDM) $N$-body simulations in the collapse phase. Especially we aim for a better understanding of the evolution of satellite halos.
Methods. To do so, we run SIDM $N$-body simulations of a low mass halo in isolation and within an external gravitational potential. The simulation setup is motivated by the perturber of the stellar stream GD-1.
Results. We find that the halo evolution is very sensitive to energy conservation errors, and a too large SIDM kernel size can artificially speed up the collapse. Moreover, we demonstrate that the King model can describe the density profile at small radii for the late stages that we have simulated. Furthermore, for our highest-resolved simulation ($N=5\times 10^{7}$) we make the data public. It can serve as a benchmark.
Conclusions. Overall, we find that the current numerical methods do not suffer from convergence problems in the late collapse phase and provide guidance on how to choose numerical parameters, e.g. that the energy conservation error is better kept well below 1%. This allows to run simulations of halos becoming concentrated enough to explain observations of GD-1 like stellar streams or strong gravitational lensing systems.
@@ 1 Introduction
A plethora of observations including the clustering of galaxies and their rotation curves can be explained by assuming an invisible matter component, dark matter (DM). Within the cosmological standard model (Lambda Cold Dark Matter, $\Lambda$CDM), this matter component mainly interacts gravitationally. Numerous experiments have been conducted to discover DM via non-gravitational interactions with standard model particles. However, unambiguous evidence for such interactions is missing (e.g. Cirelli et al., 2024). Nevertheless, there exist hints from astronomical observations that DM may have self-interactions, altering the evolution and structure of DM halos.
Self-interacting dark matter (SIDM) has historically been introduced to solve problems on galactic scales. In particular, it was meant to reduce the DM densities in the inner region of halos and to lower the abundance of satellites, i.e. suppress substructure (Spergel & Steinhardt, 2000). Even if those problems can be solved with baryonic physics and improved comparison to observations (for reviews on small-scale problems of $\Lambda$CDM see Bullock & Boylan-Kolchin, 2017; Sales et al., 2022), the underlying problem of understanding the nature of DM remains and poses one of the largest challenges in contemporary physics making it imperative to search for signatures of DM physics.
@@
There have been extensive studies on the impact of DM self-interactions on various astrophysical systems has been studied. This for example includes merging galaxy clusters (e.g. Robertson et al., 2017; Fischer et al., 2023; Sabarish et al., 2024; Valdarnini, 2024) or dwarf galaxies (e.g. Vogelsberger et al., 2014; Fry et al., 2015; Ren et al., 2019; Ebisu et al., 2022; Mancera Piña et al., 2024; Kong & Yu, 2025). Moreover, these systems have been used to constrain the strength of DM scattering as a function of the relative velocity of the particles (Kaplinghat et al., 2016; Yang et al., 2024). For a comprehensive overview, we refer to the review articles by Tulin & Yu (2018); Adhikari et al. (2022).
@@
The impact of DM self-interactions on the evolution of a DM halo can be described with the help of heat conduction (Balberg et al., 2002). The scatterings among the DM particles lead to an energy transfer, which follows the velocity dispersion gradient of the DM. This implies that a DM halo may experience heat conduction inwards to its centre when a positive velocity dispersion gradient is present. This is for example the case for an Navarro–Frenk–White (NFW) profile (Navarro et al., 1996). As a consequence of the heat flow the central density decreases.
@@
When the heat conduction continues, the velocity dispersion gradient becomes negative at all radii. This means that heat is only flowing outwards. While the central region of the halo is losing energy, it becomes more compact and the density increases. This phase of the halo evolution has been named ‘collapse phase’. Here the self-interactions lead to a runaway process resulting in the gravothermal catastrophe (Lynden-Bell & Wood, 1968; Burkert, 2000).
@@
Interestingly, some observations point towards more compact substructures, which could potentially be explained by the gravothermal collapse of SIDM halos (e.g. Turner et al., 2021; Yang & Yu, 2021; Yang et al., 2023a; Gad-Nasr et al., 2024; Dutra et al., 2024; Ragagnin et al., 2024; Shah & Adhikari, 2024; Zeng et al., 2025). This includes gravitational lensing (e.g. Meneghetti et al., 2020; Granata et al., 2022), in particular the observation of satellites acting as strong lens perturbers (e.g. Vegetti et al., 2010; Enzi et al., 2025; Despali et al., 2024; Cao et al., 2025; Li et al., 2025; Minor, 2025). In particular, Nadler et al. (2023) found that the projected logarithmic density profile slope of their SIDM simulations for the collapse phase is consistent with the observations by Minor et al. (2021). While gravitational lensing is very powerful in detecting substructures (e.g. Gilman et al., 2021; Keeley et al., 2024; Gannon et al., 2025), especially if they are dark, there are other possibilities to learn about substructures.
@@
One of them is via their impact on stellar streams (e.g. Lu et al., 2025). For example, the stellar stream GD-1, being one of the coldest and longest streams of the Milky Way, has a rich morphology (e.g. Grillmair & Dionatos, 2006; Tavangar & Price-Whelan, 2025). It shows gaps (e.g. Carlberg & Grillmair, 2013; de Boer et al., 2018, 2020; Banik et al., 2021; Malhan et al., 2022) and spur (e.g. Price-Whelan & Bonaca, 2018; Bonaca & Hogg, 2018; Bonaca et al., 2020). These structures can be explained by compact perturbers passing by the stream. Depending on the DM model we expect compact dark substructures to be more or less abundant. Zhang et al. (2025) showed that SIDM can form objects compact enough to explain the gaps and spur-like feature in GD-1, while CDM might be able to only explain the gaps.
@@
The abundance and compactness of the substructures leave their distinct impact on stellar streams. Predictions for GD-1-like streams in the scenario of collisionless DM have been derived by Adams et al. (2024). Additionally, self-interactions could directly impact the satellite from which the stellar stream originates. This is particularly relevant if the satellite is a dwarf galaxy in contrast to a star cluster. The self-interaction of the satellite’s DM halo can potentially have at least a small impact on the morphology of the stellar stream (e.g. Zeng et al., 2024; Hainje et al., 2025). For the future, we expect the wealth of observational data on stellar streams to increase strongly, not only for streams of the Milky Way (Bonaca & Price-Whelan, 2025) but also for extragalactic stellar streams (e.g. Pearson et al., 2022). This opens up an opportunity to constrain DM models, especially SIDM and eventually find evidence for non-gravitational DM interactions.
@@
Modelling DM substructures from first principles, in particular gravothermally collapsing SIDM satellites, is challenging (e.g. Yang & Yu, 2022; Zhong et al., 2023; Mace et al., 2024; Palubski et al., 2024). Problems leading to energy non-conservation in $N$-body simulations of the late collapse phase of isolated halos have been explained in detail by Fischer et al. (2024a). However, a small error in the energy conservation does not necessarily imply that the simulation results are accurate. Checking the conservation of total energy can only serve as a diagnostic, but it does not ensure reliable results. Instead, it is necessary to run a simulation with a high resolution and parameters tuned to achieve highly accurate results to obtain a benchmark against which other simulations can be tested.
@@
Several aspects concerning the accuracy of simulations of the gravothermal collapse, such as their convergence behaviour, have not yet been fully investigated. Therefore, in this paper, we study the numerical properties of such SIDM simulations using the $N$-body code OpenGadget3. In addition, we study qualitative differences arising from the velocity and angular dependence of the SIDM cross-section for a halo in isolation and as a satellite. For the latter one, our set-up is motivated by the GD-1 perturber. As a result, we comment on how to choose numerical parameters for SIDM simulations. Moreover, we show that continuing an SIDM simulation with a too large time step may lead to wrong predictions on the enclosed mass. Finally, we demonstrate that a King model provides a good description of the density profile at small radii in the collapse phase and our results can be applied to study strong lensing perturbers.
@@
The remainder of the paper is structured as follows. In Sect. 2 we describe the numerical set-up of this study. The results of isolated halo simulations are presented in Sect. 3, followed by Sect. 4 with a presentation of the findings when the halo is evolved in an external potential. In Sect. 5, we discuss the limitations of our simulations as well as directions to continue this research. Finally, we summarise and conclude in Sect. 6.
@@ 2 Numerical set-up
In this section we explain the employed numerical set-up. We begin with the simulation code and continue with the initial conditions (IC) and finally summarize the simulation parameters.
For our simulations we use the cosmological $N$-body code OpenGadget3. It is a successor of the Gadget-2 code (Springel, 2005). The domain decomposition and the neighbour search we use have been described by Ragagnin et al. (2016). Additional information on the code is also given by Groth et al. (2023) and the references therein.
OpenGadget3 contains a module to simulate DM self-interactions, introduced by Fischer et al. (2021a, b, 2022, 2024b) and has been used for numerous studies of SIDM. A unique feature of the SIDM module is that it is capable of simulating very anisotropic cross-sections (see also Fischer et al., 2021a; Arido et al., 2025). Moreover, it allows to perform the computations in parallel employing a message passing interface (MPI) parallelisation and/or an open multi-processing (OpenMP) parallelisation. At the same time it explicitly conserves energy and linear momentum. A separate time-step criterion has been implemented for the self-interactions ensuring small-enough time steps. For computing the interactions between the numerical particles, a spline kernel is employed (Monaghan & Lattanzio, 1985). The kernel size for each particle is chosen adaptively and set by the next $N_{\mathrm{ngb}}=48$ neighbours if not stated otherwise.
We choose the ICs for our simulations motivated by the perturber of the stellar stream GD-1. It should be a halo that is a reasonable progenitor evolving to a state potentially explaining the observations. In this sense, a high concentration is desirable as it makes a fast gravothermal evolution due to SIDM more plausible. Our ICs follow an NFW profile (Navarro et al., 1996),
@@
We set the density parameter to $\rho_{0}=4.42\times 10^{7}\,\mathrm{M_{\odot}}\,\mathrm{kpc}^{-3}$ and the scale radius to $r_{\mathrm{s}}=1.28\,\mathrm{kpc}$. The halo has a mass of about $\approx 2.8\times 10^{9}\,\mathrm{M_{\odot}}$ and its concentration parameter is $c\approx 22$, which is a little higher than the typically concentration at this halo mass (Dutton & Macciò, 2014). For our simulations, we generate the ICs using SpherIC (Garrison-Kimmel et al., 2013) and sample the halo up to a radius $r_{\mathrm{cut}}=10\,r_{\mathrm{s}}$ or $r_{\mathrm{cut}}=15\,r_{\mathrm{s}}$. The ICs are all without baryons, i.e. DM only.
We set a spherical boundary condition at 200 kpc to avoid problems arising from particles travelling very far from the halo. The boundary condition reflects particles that are travelling outwards, i.e. it does not directly affect energy conservation (as done by Fischer et al., 2024a).
Moreover, we simulate the halo with three different mass resolutions and also vary the gravitational softening length. The highest resolved simulations contain $N=5\times 10^{7}$ particles. Furthermore, we employ different values for the gravitational time step parameter $\eta$ (eq. 34 by Springel, 2005) or the SIDM time step parameter $\tau$ (eq. 5 by Fischer et al., 2024b) and employ for some simulations a minimum allowed time step $\Delta t_{\mathrm{min}}$. In addition, we set the opening angle for the gravitational force computations by setting the value for the parameter $\alpha=5\times 10^{-4}$ (eq. 18 by Springel, 2005). An overview with the details of our simulations is provided in Tab. 1.
For most of our simulations we assume an isotropic velocity-independent cross-section. Simulations with a different angular or velocity dependence are described in Sect. 3.8. In addition to the isolated halo simulations, we place the halo in an external potential. This allows for studying its evolution when being subject to tidal forces. The details are described in Sect. 4.1. An overview of the 29 simulations that we show in this paper is given in Appendix A. The choice for various numerical parameters, such as the softening length or the values for the time-step criteria are highlighted.
@@ 3 Isolated halo evolution
In this section, we present our simulation results and discuss the relevance of various numerical parameters for achieving accurate simulations. This includes various aspects such as the role of variable time steps, the radius up to which the ICs are sampled, the role of mass resolution, the gravitational softening length and the SIDM kernel size, as well as setting a minimal time step. Moreover, we also study different angular and velocity dependencies and comment on their qualitative differences. Furthermore, we explore the evolution of the velocity anisotropy. In addition, we study how much the evolution of an SIDM halo simulation depends on the choice for the gravitational time step criterion in Appendix B. We also demonstrate in Appendix D that the inner region during the collapse phase can be well described by a King model (King, 1962).
For all simulations, we use the peak find algorithm introduced by Fischer et al. (2021b) to find the centre of the halo. It corresponds to the minimum of the gravitational potential. We compute physical quantities such as the enclosed mass with respect to this centre.

### v4

@@ Abstract
Context. The properties of satellite halos provide a promising probe for dark matter (DM) physics. Observations have motivated current efforts to explain surprisingly compact DM halos. If DM is not collisionless, but has strong self-interactions, halos can undergo gravothermal collapse, leading to higher densities in the central region of the halo. However, it is challenging to model this collapse phase from first principles.
Aims. To improve on this, we sought to better understand the numerical challenges and convergence properties of self-interacting dark matter (SIDM) $N$-body simulations in the collapse phase. Especially, our aim was to better understand the evolution of satellite halos.
Methods. To do so, we ran SIDM $N$-body simulations of a low-mass halo in isolation and within an external gravitational potential. The simulation set-up was motivated by the perturber of the stellar stream GD-1.
Results. We find that the halo evolution is very sensitive to energy conservation errors, and a SIDM kernel size that is too large can artificially speed up the collapse. Moreover, we demonstrate that the King model can describe the density profile at small radii for the late stages that we have simulated. Furthermore, for our most highly resolved simulation ($N=5\times 10^{7}$) we have made the data public. It can serve as a benchmark.
Conclusions. Overall, we find that the current numerical methods do not suffer from convergence problems in the late collapse phase and provide guidance on how to choose numerical parameters, for example that the energy conservation error is better kept well below 1%. This allows simulations to be run of halos that become concentrated enough to explain observations of GD-1-like stellar streams or strong gravitational lensing systems.
@@ 1 Introduction
[…]
@@
There have been extensive studies on the impact of DM self-interactions on various astrophysical systems. This, for example, includes merging galaxy clusters (e.g. Robertson et al., 2017; Fischer et al., 2023; Sabarish et al., 2024; Valdarnini, 2024) or dwarf galaxies (e.g. Vogelsberger et al., 2014; Fry et al., 2015; Ren et al., 2019; Ebisu et al., 2022; Mancera Piña et al., 2024; Kong & Yu, 2025). Moreover, these systems have been used to constrain the strength of DM scattering as a function of the relative velocity of the particles (Kaplinghat et al., 2016; Yang et al., 2024). For a comprehensive overview, we refer to the review articles by Tulin & Yu (2018); Adhikari et al. (2022).
@@
The impact of DM self-interactions on the evolution of a DM halo can be described with the help of heat conduction (Balberg et al., 2002). The scatterings among the DM particles lead to an energy transfer that follows the velocity dispersion gradient of the DM. This implies that a DM halo may experience heat conduction inwards to its centre when a positive velocity dispersion gradient is present. This is, for example, the case for a Navarro–Frenk–White (NFW) profile (Navarro et al., 1996). As a consequence of the heat flow, the central density decreases.
@@
When heat conduction continues, the velocity dispersion gradient becomes negative at all radii. This means that heat only flows outwards. While the central region of the halo is losing energy, it becomes more compact and the density increases. This phase of the halo evolution has been named the collapse phase. Here, the self-interactions lead to a runaway process resulting in gravothermal catastrophe (Lynden-Bell & Wood, 1968; Burkert, 2000).
@@
Interestingly, some observations point towards more compact substructures, which could potentially be explained by the gravothermal collapse of SIDM halos (e.g. Turner et al., 2021; Yang & Yu, 2021; Yang et al., 2023a; Gad-Nasr et al., 2024; Dutra et al., 2024; Ragagnin et al., 2024; Shah & Adhikari, 2024; Zeng et al., 2025b). This includes gravitational lensing (e.g. Meneghetti et al., 2020; Granata et al., 2022), in particular the observation of satellites acting as strong lens perturbers (e.g. Vegetti et al., 2010; Enzi et al., 2025; Despali et al., 2025; Cao et al., 2025; Li et al., 2025; Minor, 2025). In particular, Nadler et al. (2023) found that the projected logarithmic density profile slope of their SIDM simulations for the collapse phase is consistent with the observations by Minor et al. (2021). While gravitational lensing is very powerful in detecting substructures (e.g. Gilman et al., 2021; Keeley et al., 2024; Gannon et al., 2025), especially if they are dark, there are other possibilities to learn about substructures.
@@
One of them is via their impact on stellar streams (e.g. Lu et al., 2025). For example, the stellar stream GD-1, which is one of the coldest and longest streams of the Milky Way, has a rich morphology (e.g. Grillmair & Dionatos, 2006; Tavangar & Price-Whelan, 2025). It shows gaps (e.g. Carlberg & Grillmair, 2013; de Boer et al., 2018; de Boer et al., 2020; Banik et al., 2021; Malhan et al., 2022) and a spur (e.g. Price-Whelan & Bonaca, 2018; Bonaca & Hogg, 2018; Bonaca et al., 2020). These structures can be explained by compact perturbers passing by the stream. Depending on the DM model, we expect compact dark substructures to be more or less abundant. Zhang et al. (2025) showed that SIDM can form objects compact enough to explain the gaps and spur-like feature in GD-1, while CDM might be able to only explain the gaps.
@@
The abundance and compactness of the substructures leave their distinct impact on stellar streams. Predictions for GD-1-like streams in the scenario of collisionless DM have been derived by Adams et al. (2025). Additionally, self-interactions could directly impact the satellite from which the stellar stream originates. This is particularly relevant if the satellite is a dwarf galaxy in contrast to a star cluster. The self-interaction of the satellite’s DM halo can potentially have at least a small impact on the morphology of the stellar stream (e.g. Zeng et al., 2025a; Hainje et al., 2025). For the future, we expect the wealth of observational data on stellar streams to increase strongly, not only for streams of the Milky Way (Bonaca & Price-Whelan, 2025), but also for extragalactic stellar streams (e.g. Pearson et al., 2022). This opens up an opportunity to constrain DM models, especially SIDM and eventually find evidence for non-gravitational DM interactions.
@@
Modelling DM substructures from first principles, in particular gravothermally collapsing SIDM satellites, is challenging (e.g. Yang & Yu, 2022; Zhong et al., 2023; Mace et al., 2024; Palubski et al., 2024). Problems leading to energy non-conservation in $N$-body simulations of the late collapse phase of isolated halos have been explained in detail in Fischer et al. (2024a). However, a small error in energy conservation does not necessarily imply that the simulation results are accurate. Checking the conservation of total energy can only serve as a diagnostic, but it does not guarantee reliable results. Instead, it is necessary to run a simulation with high resolution and parameters tuned to achieve highly accurate results to obtain a benchmark against which other simulations can be tested.
@@
Several aspects concerning the accuracy of simulations of the gravothermal collapse, such as their convergence behaviour, have not yet been fully investigated. Therefore, for this paper, we studied the numerical properties of such SIDM simulations using the $N$-body code OpenGadget3. In addition, we studied the qualitative differences arising from the velocity and angular dependence of the SIDM cross-section for a halo in isolation and as a satellite. For the latter, our set-up was motivated by the GD-1 perturber. As a result, we comment on how to choose numerical parameters for SIDM simulations. Moreover, we show that continuing an SIDM simulation with a too large time step may lead to incorrect predictions on the enclosed mass. Finally, we demonstrate that a King model provides a good description of the density profile at small radii in the collapse phase, and our results can be applied to study strong lensing perturbers.
@@
The remainder of the paper is structured as follows. In Sect. 2 we describe the numerical set-up of this study. The results of isolated halo simulations are presented in Sect. 3, followed by Sect. 4 with a presentation of the findings when the halo is evolved in an external potential. In Sect. 5 we discuss the limitations of our simulations as well as directions to continue this research. Finally, we summarise and conclude in Sect. 6. Additional information is provided in the appendices.
@@ 2 Numerical set-up
In this section we explain the employed numerical set-up. We begin with the simulation code and continue with the initial conditions (ICs) and finally summarise the simulation parameters.
For our simulations we used the cosmological $N$-body code OpenGadget3. It is a successor of the Gadget-2 code (Springel, 2005). The domain decomposition and the neighbour search we used have been described by Ragagnin et al. (2016). Additional information on the code is also given by Groth et al. (2023) and the references therein.
OpenGadget3 contains a module to simulate DM self-interactions, introduced by Fischer et al. (2021a, b, 2022, 2024b) and has been used for numerous studies of SIDM. A unique feature of the SIDM module is that it is capable of simulating very anisotropic cross-sections (see also Fischer et al., 2021a; Arido et al., 2025). Moreover, it allows one to perform the computations in parallel employing a message passing interface (MPI) parallelisation and/or an open multi-processing (OpenMP) parallelisation. At the same time, it explicitly conserves energy and linear momentum. A separate time step criterion was implemented for the self-interactions, which ensures small-enough time steps. For computing the interactions between numerical particles, a spline kernel was employed (Monaghan & Lattanzio, 1985). The kernel size for each particle was chosen adaptively and set by the next $N_{\mathrm{ngb}}=48$ neighbours if not stated otherwise.
We chose the ICs for our simulations motivated by the perturber of the stellar stream GD-1. It should be a halo that is a reasonable progenitor evolving to a state potentially explaining the observations. In this sense, a high concentration is desirable, as it makes rapid gravothermal evolution driven by SIDM more plausible. Our ICs follow an NFW profile (Navarro et al., 1996):
@@
We set the density parameter to $\rho_{0}=4.42\times 10^{7}\,\mathrm{M_{\odot}}\,\mathrm{kpc}^{-3}$ and the scale radius to $r_{\mathrm{s}}=1.28\,\mathrm{kpc}$. The halo has a mass of about $\approx 2.8\times 10^{9}\,\mathrm{M_{\odot}}$ and its concentration parameter is $c\approx 22$, which is slightly higher than the typical concentration at this halo mass (Dutton & Macciò, 2014). For our simulations, we generated the ICs using SpherIC (Garrison-Kimmel et al., 2013) and sampled the halo up to a radius $r_{\mathrm{cut}}=10\,r_{\mathrm{s}}$ or $r_{\mathrm{cut}}=15\,r_{\mathrm{s}}$. The ICs are all without baryons, i.e. DM only.
We set a spherical boundary condition at 200 kpc to avoid problems arising from particles travelling very far from the halo. The boundary condition reflects particles that are travelling outwards, i.e. it does not directly affect energy conservation (as done by Fischer et al., 2024a). The self-interactions locally thermalise the velocity distribution and consequently populate the high-velocity tail of the Maxwell-Boltzmann distribution. This leads to a few particles escaping the halo. When those particles travel to large radii, this can lead to significant round-off errors in the gravity calculations, given that we store the positions of the particles using double-precision floating-point numbers. In our case, the precision of the particles may not be of much concern, given that we do not run our simulations for a very long time, and thus the particles may not travel to very large distances. However, to ensure that no problem arises from a reduced precision in the position of the particles, we introduced the boundary condition.
Moreover, we simulated the halo with three different mass resolutions and also varied the gravitational softening length. The most highly resolved simulations contain $N=5\times 10^{7}$ particles. Furthermore, we employed different values for the gravitational time step parameter $\eta$ (eq. 34 by Springel, 2005) or the SIDM time step parameter $\tau$ (eq. 5 by Fischer et al., 2024b) and employed for some simulations a minimum allowed time step $\Delta t_{\mathrm{min}}$. In addition, we set the opening angle for the gravitational force computations by setting the value of the parameter $\alpha=5\times 10^{-4}$ (eq. 18 by Springel, 2005). An overview with the details of our simulations is provided in Table 1.
For most of our simulations we assumed an isotropic velocity-independent cross-section. Simulations with a different angular or velocity dependence are described in Sect. 3.8. In addition to the isolated halo simulations, we placed the halo in an external potential. This allowed us to study its evolution when being subject to tidal forces. The details are described in Sect. 4.1. An overview of the 29 simulations that we show in this paper is given in Appendix A. The choices of the various numerical parameters, such as the softening length or the values for the time step criteria, are highlighted.
@@ 3 Isolated halo evolution
In this section we present our simulation results and discuss the relevance of various numerical parameters for achieving accurate simulations. This includes various aspects such as the role of variable time steps, the radius up to which the ICs are sampled, the role of mass resolution, the gravitational softening length and the SIDM kernel size, as well as setting a minimal time step. We also present results for different angular and velocity dependences and comment on their qualitative differences. Furthermore, we explore the evolution of the velocity anisotropy. In addition, we discuss how much the evolution of an SIDM halo simulation depends on the choice for the gravitational time step criterion in Appendix B. We also demonstrate in Appendix D that the inner region during the collapse phase can be well described by a King model (King, 1962).
For all simulations, we used the peak find algorithm introduced by Fischer et al. (2021b) to find the centre of the halo. It corresponds to the minimum of the gravitational potential. We computed physical quantities such as the enclosed mass with respect to this centre.